OpenStudy (anonymous):

☻WILL GIVE MEDAL AND FAN ☻ Harrison and Sherrie are making decisions about their bank accounts. Harrison wants to deposit $200 as a principle amount, with an interest of 2% compounded quarterly. Sherrie wants to deposit $200 as the principle amount, with an interest of 4% compounded monthly. Explain which method results in more money after 2 years. Show all work.

OpenStudy (anaise):

interest rate?

OpenStudy (calculusxy):

I am that good at these kinds of math honestly, but I love to try them out! Harrison Principal (starting $) : \(\large $200\) Quarterly (every 4 months, I believe): \(\large 12 \div 4 = 3\) Interest per 3 months : \(0.02 \times 3\) \(\large $200 + (0.02 \times 3)\) Sherrie Principal (starting $) : \(\large $200\) Quarterly (every moths): \(\large 12 \) Interest per 3 months : \(0.04 \times 12\) \(\large $200 + (0.04 \times 12)\)

OpenStudy (calculusxy):

*I am NOT that good

OpenStudy (anonymous):

Lol, I just multiplied the percent over and over again

OpenStudy (anonymous):

Harrison: 200 * 0.02 + 200 = 204 204 * 0.02 + 204 = 208.08 208.08 * 0.02 + 208.08 = 212.24 212.24 * 0.02 + 212.24 = 216.49 216.49 * 0.02 + 216.49 = 220.82 220.82 * 0.02 + 220.82 = 225.24 225.24 * 0.02 + 225.24 = 229.75 229.75 * 0.02 + 229.75 = 234.35 *** 234.35 *** Sherrie: 200 * 0.04 + 200 = 208 208 * 0.04 + 208 = 216.32 216.32 * 0.04 + 216.32 = 224.97 224.97 * 0.04 + 224.97 = 233.97 233.97 * 0.04 + 233.97 = 243.33 *** Sherrie would earn more in 5 months than Harrison would in 2 years ***

OpenStudy (anonymous):

Guess there's an easier way. Thanks for your help anyways!

OpenStudy (calculusxy):

No problem :)